OpenStudy (anonymous):

(08.02)Which description best describes the solution to the following system of equations? y = −one halfx + 9 y = x + 7 Lines y = −one halfx + 9 and y = x + 7 intersect the x-axis. Lines y = −one halfx + 9 and y = x + 7 intersect the y-axis. Line y = −one halfx + 9 intersects the origin. Line y = −one halfx + 9 intersects line y = x + 7. PLEASE I NEED HELP THIS PREALGEBRA STUFF IS HARD

OpenStudy (anonymous):

HELP ME PPPPPPPPPPPLLLLLLLLLLLLLLEEEEEEEEEEEEAAAAAAAAAAAASSSSSSSSSSEEEEEEEEEEE

ganeshie8 (ganeshie8):

Solution to a system of linear equations is the intersecting point of all the lines

OpenStudy (anonymous):

So which one is it please help I need to pass

OpenStudy (usukidoll):

graph y=x+7 and \[y=\frac{1}{2}x+9 \] let x = be any number like x= -2,-2,0,1,2 and graph until the lines intersect.

OpenStudy (usukidoll):

and we're not supposed to be giving direct answers as it is against OpenStudy's Code of Conduct. We can only guide the user.

ganeshie8 (ganeshie8):